OpenStudy (anonymous):

Carl conducted an experiment to determine if the there is a difference in mean body temperature between men and women. He found that the mean body temperature for men in the sample was 91.1 with a population standard deviation of 0.52 and the mean body temperature for women in the sample was 97.6 with a population standard deviation of 0.45.

OpenStudy (anonymous):

Assuming the population of body temperatures for men and women is normally distributed, calculate the 98% confidence interval and the margin of error for the mean body temperature for both men and women. Using complete sentences, explain what these confidence intervals mean in the context of the problem.

OpenStudy (anonymous):

@PlsHaveMercyLilB @SolomonZelman

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@IMStuck do you think you can help

OpenStudy (imstuck):

I am trying really hard...

OpenStudy (anonymous):

Same number of men as there are women?

OpenStudy (anonymous):

I would assume so, judging by the fact that that info isn't listed. The confidence interval will have the form \[\left((\bar{x}_1-\bar{x}_2)-Z_{\alpha/2}\sqrt{\frac{{\sigma_1}^2}{{n_1}^2}+\frac{{\sigma_2}^2}{{n_2}^2}},~(\bar{x}_1-\bar{x}_2)+Z_{\alpha/2}\sqrt{\frac{{\sigma_1}^2}{{n_1}^2}+\frac{{\sigma_2}^2}{{n_2}^2}}\right)\] where \(\bar{x}_1-\bar{x}_2\) is the estimated mean difference, \(Z_{\alpha/2}\) is the critical \(z\) value for a \((1-\alpha)100\%\) confidence level, \(\sigma_1\) and \(\sigma_2\) are the corresponding standard deviations, and \(n_1\) and \(n_2\) are the respective sample sizes. You have all the info necessary to construct the interval except for the sample sizes.

OpenStudy (anonymous):

if have to do all that? @SithsAndGiggles

OpenStudy (anonymous):

Well yes, I'm not planning on doing it for you. Do you have the sample size(s) given?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

That's strange. You need the info to proceed with the interval :(

OpenStudy (anonymous):

everything i listed at the top is given

OpenStudy (anonymous):

@abb0t @aaronq @aum @AriPotta can yall help me

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@AriPotta u think u can help

OpenStudy (aripotta):

maybe. i'm thinking

OpenStudy (anonymous):

please help me i am so lost... i really dont know how to do this!

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

Well, I've done what I can... Is this problem taken from a textbook, written by a professor, etc?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can anybody tell me what the confidence intervals @campbell_st @ash2326 @ash2326 @abb0t @AriPotta

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

sorry, i can't help. i keep losing my internet connection

OpenStudy (anonymous):

awwwww

OpenStudy (anonymous):

@Prestianne help

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@abb0t help please

OpenStudy (anonymous):

hey, what's wrong??

OpenStudy (anonymous):

i was looking for some answers to ur first question and i found something, does this help?? Make sure you read the whole thing, its not that long when u read it carefully. go to that website: http://openstudy.com/study#/updates/5272cc39e4b077aaee848854

OpenStudy (anonymous):

Looks like those guys ran into the same problem - no sample size is given. As for the margin of error - it's the \(Z_{\alpha/2}\sqrt{\cdots}\) term. Not having the sample size is the issue, though. You simply can't answer this question without it.